OpenStudy (anonymous):

MEDALS!!!! Please help solve problem. Express answer as an integer or simplified fraction. x/6-4=x/3-3

OpenStudy (anonymous):

move x/3 to the left side x/6-x/3 -4=-3 add 4 to both sides x/6-x/3=1 \[\frac{ x-2x }{ 6}=1\] -x/6=1 x=-6

OpenStudy (anonymous):

@JANELLSM98